OpenStudy (anonymous):

If two number cubes are rolled at the same time what is the probability that the two cubes show different numbers?

OpenStudy (anonymous):

Someone help!

Parth (parthkohli):

Okay, so we'd multiply the probabilities here.

Parth (parthkohli):

In the first cube, there can be 6 possibilities outta 6, but in the second one, there'd be 5 outta 6 because no number can repeat.

Parth (parthkohli):

So you gotta multiply 6/6 and 5/6 and that'd be your complete answer.

OpenStudy (anonymous):

Wow, I totally over thought that.

Parth (parthkohli):

lol

OpenStudy (anonymous):

THANK YOU!

Parth (parthkohli):

Y0U'R3 W31C0M3!

OpenStudy (anonymous):

i've got another one for you, :3 hahah. A DJ for a school dance has a CD with 6 slow songs and 5 fast songs on it. As he plays each song he removes it from the playlist. What is the probability that the first two songs he plays are slow?

Parth (parthkohli):

Oh this one is easy again haha There are total 11 songs The first one has a probability of 6/11 The second one then has a probability of 5/10 because one CD is removed. Now just multiply 6/11 and 5/10 =)

OpenStudy (anonymous):

there are 36 possible cases out of them number of cases having same number on cubes are 1,1 2,2 3,3 4,4 5,5 and 6,6 this means that number of cases having diff number is 30 prob = 30/36=5/6

Parth (parthkohli):

All I can say is that is not correct ^ :|

OpenStudy (anonymous):

the ques says both are rolled at same time

Parth (parthkohli):

Oh lol I was talking about the second one..

OpenStudy (anonymous):

ohk.. i just chk we both got same ans for 1 ques..i am confused y r u saying this..:/

Parth (parthkohli):

lol yes you were correct in the first one

OpenStudy (anonymous):

How about (2k+4)^4

Parth (parthkohli):

That's binomial theorem. Can you close this question and post that in another one please?

OpenStudy (anonymous):

Yeah.